Grade 11General Physics

A helicopter which is flying 60* north of east lands itself 45* east of north due to the wind.The speed of the helicopter is 100mile/h relative to air. Calculate to speed of the wind if it is blowing from west to east

JAYPRAKASH MAURYA
9 Years agoGrade 11

1 Answer

Vikas TU
ApprovedApproved Tutor Answer9 Years ago
Dear Student,
The speed of the helicopter is 100mile/h with respect to air. 
 
Give v1 a chance to be the speed of the helicopter 
 
What's more, v2 be the speed of the breeze which has greatness b 
 
v1 = 100 cos 60o i + 100 sin 60o j 
 
= 50i + 50√3j 
 
What's more, from the other idea 
 
v2 = an i 
 
the resultant is given by 
 
v1 = v1+v2 
 
=(50+a)i + 50√3j 
 
As given in the inquiry , the helicopter lands 45* east of north because of the breeze, at that point x part will be equivalent to y segment 
 
50+a = 50√3 
 
a = 50(√3-1) 
 
a = 36.6 miles/hr 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)